# Optimal ambition in business, politics and life

### v3

@@ Abstract
In business, politics and life, folk wisdom encourages people to aim for above-average results, but to not let the perfect be the enemy of the good. Here, we mathematically formalize and extend this folk wisdom. We model a time-limited search for strategies having uncertain rewards. At each time step, the searcher either is satisfied with their current reward or continues searching. We prove that the optimal satisfaction threshold is both finite and strictly larger than the mean of available rewards—matching the folk wisdom. We show that being too ambitious has a higher expected cost than being too cautious. We show that the optimal satisfaction threshold increases if the search time is longer, or if the reward distribution is rugged (i.e., has low autocorrelation) or left skewed. The skewness result reveals counterintuitive contrasts between optimal ambition and optimal risk taking. We show that using upward social comparison to assess the reward landscape substantially harms expected performance. We show how these insights can be applied in real-world settings, using examples from entrepreneurship, economic policy, political campaigns, online dating and college admissions.
@@
Norman Vincent Peale famously said, “Shoot for the Moon. Even if you miss, you’ll land among the stars.” However, folk wisdom suggests that it is actually possible to aim too high, just as it is possible to aim too low. In a business context, entrepreneurs are encouraged to pursue big ideas, but not unrealistic pie-in-the-sky ones. Career advisers encourage people to build skills and social capital gradually, rather than expecting their first job to be perfect (e.g., ref. Jay (2012)), and to be opportunistically mobile but not restless or non-committal. Common dating advice encourages people to have realistic expectations about what to look for in a partner. In politics, incrementalism encourages gradual changes to improve society without destabilizing its institutions Lindblom (1959). Dr. Martin Luther King Jr. famously said that “the arc of the moral universe is long, but it bends toward justice” King (March 31, 1968).
@@
Empirical research seems to support these folk intuitions encouraging intermediate ambition. For example, intermediate procrastinators are often the most successful entrepreneurs and innovators Grant (2017). They do not over-commit to their first ideas, but they still have the focus to execute better ideas later. Employees’ propensities to change jobs tend to have mixed effects on career success and professional well-being Guan et al. (2019). Research supports incrementalism as a policymaking strategy, due to checks and balances, distributed expertise, and pluralistic social and political interests Lindblom (1959); Hayes (2002). Utopianism has sometimes produced tragic consequences such as the Great Chinese Famine in Mao Zedong’s China Courtois (1999); Pinker (2012). The happiest marriages, on average, occur between people who are similar to each other in traits related to desirability (e.g., health, vitality, exercise, spirituality) George et al. (2015). Natural resource exploiters such as fishers, who do not have access to perfect information about where the resources are most plentiful, maximize their catch by aiming for patches of slightly above-average quality Burgess et al. (2020a). These are a few examples of many.
@@
Here, we mathematically formalize and extend these folk intuitions about ambition. We show in a general search model that optimal ambition targets outcomes that are both strictly finite and strictly larger than the mean of available rewards. In other words, optimal ambition does not ‘shoot for the moon’, but it does try to do better than average. Our model has conceptual parallels to areas of applied mathematics, economics, and theoretical biology related to optimal search and optimal stopping, which we discuss. We show how optimal ambition depends on certain features of the reward landscape, and how optimal ambition is affected by social comparison. Our results provide a precise and accessible conceptual link between the diverse domains in which intermediate ambition seems to be supported. Our model provides actionable insights for optimally calibrating ambition, which we illustrate using empirical examples from economic policy, entrepreneurship, and political campaigns. Our model also provides testable hypotheses for empirical research, which we illustrate using examples from online dating and college admissions.
I Model
@@
Strategies and rewards could abstractly represent, for example, jobs and salaries for a job-seeker, colleges and application cycles for a prospective student, venture ideas and payoffs for an entrepreneur, dating prospects and relationship satisfaction for a single, campaign strategies and popularity for a political candidate, policies and their outcomes for a policy maker, or patches with different potential harvests for a harvester. There are many other possible examples. Our model is admittedly simpler than these real decision settings, but we argue that its key insights should generalize to more complex settings.
@@
The agent must balance the costs of being too easily satisfied and settling for less than what is attainable, with the costs of being too ambitious and passing on high, attainable rewards while searching fruitlessly for the unattainable. We prove that the optimal satisfaction threshold, $T$, is strictly larger than the mean, $\mu$ (i.e., agents should try to do better than average), but the optimal $T$ is also finite (i.e., overambition is possible, and costly). We explore how the optimal satisfaction threshold depends on several properties of the search, the reward distribution, and the information agents use to make their decisions.
@@
Figure 1: Stylized representations of landscape ruggedness (left) and skewness (right). Ruggedness determines how different successive rewards are from each other. Skewness determines the relative abundance of low and high rewards, even for landscapes that share mean and variance.
The agent’s reward in time period $t$, $X_{t}$, is given by:
@@
Here, $\epsilon_{t}$ are independent and identically distributed (i.i.d.) samples from the reward distribution. Our model is effectively an autoregressive (AR1) random walk, which ends at time $t_{\text{max}}$ ($t\leq t_{\text{max}}<\infty$). This approach has the advantage of model parsimony while capturing the properties of commonly used fitness landscapes (e.g., the underlying distribution Skellett et al. (2005) and parent–offspring correlation Verel et al. (2003) in fitness landscapes, and autocorrelation behavior on general neighborhood digraph landscapes Dimova et al. (2005)). Equation (1) describes the two phases of the search. In the ‘explore’ phase, each agent conducts a random walk. Once the agent meets their respective threshold, they initiate the ‘exploit’ phase and remain at the same reward value until they run out of time.
@@
Our model offers a general, tractable, and accessible formulation applicable to a widely studied class of sequential decision problems that aim to maximize rewards (or minimize costs) under uncertainty (e.g. MacQueen and Miller Jr (1960); van Moerbeke (1974); Freeman (1983); Powell (2019); Kohn and Shavell (1974)). Optimal stopping problems, concerned with determining the best time to take a particular action, have been widely applied in management science Ciocan and Mišić (2022) and finance Liu and Mu (2022). Optimal search has been explored in economics Kohn and Shavell (1974), and in theoretical biology via optimal foraging theory Pyke (1984). For example, the Marginal Value Theorem Charnov (1976) has been used to explain how long diving mammals spend underwater Kramer (1988), how long birds spend in any given tree Cowie (1977), and how long bumblebees spend on flowers Schmid-Hempel et al. (1985). These same ideas have even helped explain the evolution of adaptations Parker and Smith (1990). Depending on the context, solving such optimal stopping or optimal search problems may require advanced computational methods Becker et al. (2019); Ciocan and Mišić (2022). In contrast, in this work, we focus on a simple, self-contained formulation that emphasizes parsimony. For a detailed mathematical treatment of optimal stopping and hitting times in autoregressive processes, see refs.  Novikov (2009); Christensen (2012); Christensen et al. (2011). We extend our model to incorporate features that are psychologically relevant and commonly reflected in human decision-making, such as varying reward distribution propertiesGuan et al. (2014) and social comparison Guan et al. (2014); Gerber et al. (2018); Muller and Fayant (2010).
II Results
II.1 Expected rewards vs. satisfaction threshold
@@ II Results
We can analytically derive the relationship between an agent’s satisfaction threshold and their expected reward, in our model, in the special case of a maximally rugged landscape ($\varphi=0$) and a Gaussian distribution of rewards, $\epsilon_{t}\sim\mathcal{N}(\mu=0,\sigma^{2})$ (see Methods). In other cases, we can accurately compute the relationship numerically (Fig. 2a).
@@
Figure 2: Panel a: analytical and numerical calculations of the expected reward distribution, as a function of the target threshold, on a maximally rugged landscape. The expected reward distribution is unimodal and negatively skewed, with an optimal threshold above zero (the landscape mean). The analytical expression given by Equation (2) (black, dashed) matches the simulation results (red). Parameter values are: $\epsilon_{t}\sim\mathcal{N}(0,1)$, $\varphi=0$, and $t_{\text{max}}=1000$. Results shown are averaged over $10^{4}$ simulations. We note that many real-world applications occur on shorter time scales ($t_{\text{max}}\ll 1000$), and consequently have lower optimal thresholds (e.g., optimal $T=1.6$ with $t_{\text{max}}=100$). Panels b and c: examples of real-world search strategies. Panel b: When online dating, heterosexual men (blue) and women (red) are most likely to message potential partners who are slightly more desirable than they are (averaged over the four cities shown in fig. 2 of ref. Bruch and Newman (2018)). Panel c: When applying for college, members of the 2008 U.S. high-school graduating class concentrated their applications on schools with median SAT scores similar to, or slightly below, their own score (data from ref. Hoxby and Avery (2012)). This suggests students were either being sub-optimally ambitious, or they faced other constraints (e.g., on income or geography), as ref. Hoxby and Avery (2012) suggests was the case.
@@
In the Methods section and Appendix A, we prove that the optimal threshold ($T$) is strictly greater than the mean reward ($\mu$). The proof is somewhat complex, but the intuition is simple. Suppose that there are two time steps ($t_{\text{max}}=2$). Searching in the second time step will yield an expected reward of $\mu$. Settling for anything less than $\mu$ after the first time step would therefore result in a lesser expected reward in the second time step.
@@
We also prove that the optimal threshold ($T$) increases in the search time ($t_{\text{max}}$), and is finite as long as the search time is finite (see Methods and Appendix A). Again, both results are intuitive: more search time increases the likelihood and cumulative payoff of achieving an ambitious target.
@@
Overshooting the optimal threshold is costlier than undershooting. In other words, an agent overshooting the optimal threshold is expected to receive lower rewards than an agent undershooting the optimal threshold by the same amount (Figs. 2a, 3, and 4). This implies that uncertainty over the reward distribution should reduce the optimal satisfaction threshold (i.e., it motivates caution), analogous to the way that left-skewed uncertainty motivates risk aversion in other contexts Benuzzi and Ploner (2024). However, in settings where agents learn about the reward distribution through searching and they can adjust their satisfaction thresholds accordingly, Kohn and Shavell Kohn and Shavell (1974) showed that uncertainty can increase the optimal satisfaction threshold due to the information value of searching.
@@
Fig. 2b and c illustrate two real-world contexts in which people seem to behave as though they understand the search tradeoff that our theory captures. Fig. 2b compares the frequencies with which heterosexual men and women in four U.S. cities sent messages to prospective partners on a dating app, as a function of their prospective partners’ desirability compared to their own desirability (data from ref. Bruch and Newman (2018)). Long-term partners tend to be similar in their desirability George et al. (2015). Thus, a zero desirability gap might be analogous to the mean reward ($\mu$) in our model. In the online dating market, both men and women expend the greatest (modal) search effort on prospective partners slightly more desirable than themselves, as our theory would predict (Fig. 2b).
@@
Fig. 2c shows the distribution of college applications from the 2008 U.S. high-school graduating class, comparing students’ own scores on the Scholastic Aptitude Test (SAT) to those of the median current student at the college they were applying to (data from ref. Hoxby and Avery (2012)). Attending a college where one has the median SAT score might be analogous to the mean reward ($\mu$) in our model. As in the dating example, the modal search (application) effort occurred near this mean college. However, unlike the dating example, students sent more applications to weaker colleges than they sent to stronger ones. Hoxby Hoxby and Avery (2012) showed that most applications in this left tail came from low- and middle-income students, who faced greater non-academic constraints (e.g., geographic, economic) on which colleges they could attend. This low-ambition tail in college applications may also relate to the ‘safety school’ strategy, which prefers unambitious options in college to striking out and being left with no options. College applicants are able to pursue multiple strategies (applications) simultaneously (albeit with some costs from time and application fees), unlike the agents in our model.
@@
II.2 Reward landscape left skewness and ruggedness increase optimal satisfaction thresholds
@@
Rugged landscapes create higher-variance rewards, which increases achievable expected rewards and optimal satisfaction thresholds (Fig. 3a). Thus, agents should be more ambitious on rugged landscapes. The reward distribution for the first time period is the same for smooth and rugged landscapes, but autocorrelation reduces the subsequent variance on smooth landscapes (see Methods). This means that agents can only make incremental changes: setting their sights too high will likely result in never encountering the desired reward, which would require a gradual climb due to the aurocorrelation.
@@
Figure 3: Smoother reward landscapes yield lower cumulative rewards than the maximally rugged, non-skewed landscapes (a). Left-skewed reward landscapes yield higher cumulative rewards (b). Smooth landscapes are generated analytically (see Methods). For the skewed landscape simulations, $\epsilon_{t}\sim\mathcal{SN}(0,1,\alpha)$, $\varphi=0$, $t_{\text{max}}=1000$, where the parameter $\alpha$ is varied to result in skew values from $-0.8$ to $0.8$, each averaged over $10^{4}$ simulations.
II.3 Social comparison hurts performance and penalizes ambition
We have previously assumed that agents know the reward distribution. What if they instead estimate the distribution, and evaluate their satisfaction, by looking at peers’ rewards Goethals and Darley (1977)? This type of behavior is called social comparison Suls et al. (2002); Schulz (2015); Lévy-Garboua and Montmarquette (2004); Bygren (2004). Indeed, in reality people often focus specifically on peers receiving higher rewards than they are—‘upward social comparison’ Gerber et al. (2018); Muller and Fayant (2010).

### v4

@@ Abstract
In business, politics and life, folk wisdom encourages people to aim for above-average results, but to not let the perfect be the enemy of the good. Here, we mathematically formalize and extend this folk wisdom. We model a time-limited search for strategies having uncertain rewards. At each time step, the searcher either is satisfied with their current reward or continues searching. We prove that the optimal satisfaction threshold is both finite and strictly larger than the mean of available rewards—matching the folk wisdom. This result is robust to search costs, unless they are high enough to prohibit all search. We show that being too ambitious has a higher expected cost than being too cautious. We show that the optimal satisfaction threshold increases if the search time is longer, or if the reward distribution is rugged (i.e., has low autocorrelation) or left-skewed. The skewness result reveals counterintuitive contrasts between optimal ambition and optimal risk taking. We show that using upward social comparison to assess the reward landscape substantially harms expected performance. We show how these insights can be applied qualitatively to real-world settings, using examples from entrepreneurship, economic policy, political campaigns, online dating and college admissions. We discuss implications of several possible extensions of our model, including intelligent search, reward landscape uncertainty and risk aversion.
I Introduction
@@
Norman Vincent Peale famously said, “Shoot for the Moon. Even if you miss, you’ll land among the stars.” However, folk wisdom suggests that it is actually possible to aim too high, just as it is possible to aim too low. In a business context, entrepreneurs are encouraged to pursue big ideas, but not unrealistic pie-in-the-sky ones. Career advisers encourage people to build skills and social capital gradually, rather than expecting their first job to be perfect (e.g., ref. [1]), and to be opportunistically mobile but not restless or noncommittal. Common dating advice encourages people to have realistic expectations about what to look for in a partner. In politics, incrementalism encourages gradual changes to improve society without destabilizing its institutions [2].
@@
Empirical research seems to support these folk intuitions encouraging intermediate ambition. For example, intermediate procrastinators are often the most successful entrepreneurs and innovators [3]. They do not over-commit to their first ideas, but they still have the focus to execute better ideas later. Employees’ propensities to change jobs tend to have mixed effects on career success and professional well-being [4]. Research supports incrementalism as a policymaking strategy, due to checks and balances, distributed expertise, and pluralistic social and political interests [2, 5]. The happiest marriages, on average, occur between people who are similar to each other in traits related to desirability (e.g., health, vitality, exercise, spirituality) [6]. Natural resource exploiters such as fishers, who do not have access to perfect information about where the resources are most plentiful, maximize their catch by aiming for patches of slightly above-average quality [7]. These are a few examples of many.
@@
Here, we mathematically formalize and extend these folk intuitions about ambition. We show in a general search model that optimal ambition targets outcomes that are both strictly finite and strictly larger than the mean of available rewards. In other words, optimal ambition does not ‘shoot for the moon’, but it does try to do better than average. Our model has conceptual parallels to areas of applied mathematics, economics and theoretical biology related to optimal search and optimal stopping, which we discuss.
We also show how optimal ambition depends on certain features of the reward landscape, and how optimal ambition is affected by social comparison and search costs. We discuss other possible extensions of our model.
Our results provide a precise and accessible conceptual link between the diverse domains in which intermediate ambition seems to be supported. Our model also provides qualitative insights for optimally calibrating ambition, which should generalize beyond our simplifying assumptions. We illustrate some of these insights using empirical examples from economic policy, entrepreneurship and political campaigns. Our model provides testable hypotheses for empirical research, which we illustrate using examples from online dating and college admissions.
II Model
@@
Strategies and rewards could abstractly represent, for example, jobs and salaries for a job-seeker, colleges and application cycles for a prospective student, venture ideas and payoffs for an entrepreneur, dating prospects and relationship satisfaction for a single, campaign strategies and popularity for a political candidate, policies and their outcomes for a policy maker, or patches with different potential harvests for a harvester. There are many other possible examples. Our model is admittedly simpler than these real decision settings, but we argue that its key qualitative insights should generalize to more complex settings.
@@
The agent must balance the costs of being too easily satisfied and settling for less than what is attainable, with the costs of being too ambitious and passing on high, attainable rewards while searching fruitlessly for the unattainable. We prove that the optimal satisfaction threshold, $T$, is strictly larger than the mean, $\mu$ (i.e., agents should try to do better than average), but the optimal $T$ is also finite (i.e., overambition is possible, and costly). We explore how the optimal satisfaction threshold depends on several properties of the search, the reward distribution, the information agents use to make their decisions and the search cost.
@@
Figure 1: Reward landscapes. Stylized representations of reward landscape ruggedness (left) and skewness (right). Ruggedness determines how different successive rewards can be from each other. Skewness determines the relative abundance of low and high rewards, among landscapes that share the same mean and variance.
With no search cost (we consider costs in section E below), the agent’s reward in time period $t$, $X_{t}$, is given by:
@@
Here, $\epsilon_{t}$ are independent and identically distributed (i.i.d.) samples from the reward distribution. Our model is effectively an autoregressive (AR(1)) random walk, which ends at time $t_{\text{max}}$ ($t\leq t_{\text{max}}<\infty$). This approach has the advantage of model parsimony while capturing the properties of commonly used fitness landscapes (e.g., the underlying distribution [8] and parent–offspring correlation [9] in fitness landscapes, and autocorrelation behavior on general neighborhood digraph landscapes [10]). Equation (1) describes the two phases of the search. In the “explore” phase, each agent conducts a random walk. Once the agent meets their respective threshold, they initiate the “exploit” phase and remain at the same reward value until they run out of time.
@@
Our model offers a general, tractable, and accessible formulation applicable to a widely studied class of sequential decision problems that aim to maximize rewards (or minimize costs) under uncertainty (e.g. [11, 12, 13, 14, 15]). Optimal stopping problems, concerned with determining the best time to take a particular action, have been widely applied in management science [16] and finance [17]. Optimal search has been explored in economics [15], and in theoretical biology via optimal foraging theory [18]. For example, the Marginal Value Theorem [19] has been used to explain how long diving mammals spend underwater [20], how long birds spend in any given tree [21] and how long bumblebees spend on flowers [22]. These same ideas have helped to explain the evolution of adaptations [23]. Depending on the context, solving such optimal stopping or optimal search problems may require advanced computational methods [24, 16]. In contrast, in this work, we focus on a simple, self-contained formulation that emphasizes parsimony. For a detailed mathematical treatment of optimal stopping and hitting times in autoregressive processes, see refs.  [25, 26, 27]. We extend our model to incorporate features that are psychologically relevant and commonly reflected in human decision-making, such as varying reward distribution properties [28], social comparison [28, 29, 30] and search costs [15].
III Results
III.1 Expected rewards vs. satisfaction threshold
@@ III Results
We can analytically derive the relationship between an agent’s satisfaction threshold and their expected reward, in our model, in the special case of a maximally rugged landscape ($\varphi=0$) and a Gaussian distribution of rewards, $\epsilon_{t}\sim\mathcal{N}(\mu=0,\sigma^{2})$. In other cases, we can accurately compute the relationship numerically (Fig. 2a).
Figure 2: Expected rewards vs. satisfaction thresholds, and empirical examples. Panel a: analytical and numerical calculations of the expected reward distribution, as a function of the target threshold, on a maximally rugged landscape. The expected reward distribution is unimodal and negatively skewed, with an optimal threshold above zero (the landscape mean). The analytical expression given by equation (2) (black, dashed) matches the simulation results (red). Parameter values are: $\epsilon_{t}\sim\mathcal{N}(0,1)$, $\varphi=0$, and $t_{\text{max}}=1000$. Results shown are averaged over $10^{4}$ simulations. We note that many real-world applications occur on shorter time scales ($t_{\text{max}}\ll 1000$), and consequently have lower optimal thresholds (e.g., optimal $T\approx 1.6$ with $t_{\text{max}}=100$). Panels b and c: examples of real-world search strategies. Panel b: When online dating, heterosexual men (blue) and women (red) are most likely to message potential partners who are slightly more desirable than they are (averaged over the four cities shown in fig. 2 of ref. [31]). Panel c: When applying for college, members of the 2008 U.S. high-school graduating class concentrated their applications on schools with median SAT scores similar to, or slightly below, their own score (data from ref. [32]). This suggests students were either being sub-optimally ambitious, or they faced other constraints (e.g., on income or geography), as ref. [32] suggests was the case.
With a rugged landscape and a Gaussian reward distribution, with probability density function (PDF) $\phi(x)$ and cumulative density function (CDF) $\Phi(x)$, we compute the expected reward by thinking of the satisfaction threshold as a divider between the ‘explore’ phase, where an agent is searching for a new strategy at each time step, and the ‘exploit’ phase, where the agent is sticking with their satisfactory strategy and its reward. The rewards collected during the ‘explore’ phase are, by definition, less than the threshold value (otherwise, the agent would stop exploring), so that $X_{\text{explore}}<T$. The rewards collected during the ‘exploit’ phase are greater than or equal to the threshold value: $X_{\text{exploit}}\geq T$. Rewards earned during the exploration phase have the lower-tail truncated normal distribution with the mean $\mu_{\text{explore}}$, and the single reward which is repeatedly collected during the exploitation phase has the upper-tail truncated normal distribution with the mean $\mu_{\text{exploit}}$.
For threshold $T$, the mean $\mu_{\text{explore}}(T)=-\dfrac{\phi(T)}{\Phi(T)}$ is given by the inverse Mills ratio [33]. The agents expect to do worse than the landscape mean in the exploration phase, on average ($\mu_{\text{explore}}<0$). The expected reward earned during the exploitation phase is given by $\mu_{\text{exploit}}(T)=\dfrac{\phi(T)}{1-\Phi(T)}>0$.
$\mathbb{E}[\text{reward}]$ $=\sum_{t_{x}=0}^{t_{\text{max}}}\Bigg[\mu_{\text{explore}}(T)t_{x}+\mu_{\text{exploit}}(T)\left(t_{\text{max}}-t_{x}\right)\Bigg]$
$\quad\times(1-\Phi(T))\Phi(T)^{t_{x}}.$ (2)
To standardize our results across different time spans ($t_{\text{max}}$), we normalize the cumulative reward by dividing by $t_{\text{max}}$ and compare expected reward per time step (e.g., Fig. 2a).
@@
In Appendix A, we prove that the optimal threshold ($T$) is strictly greater than the mean reward ($\mu$). The proof is somewhat complex, but the intuition is simple. Suppose that there are two time steps ($t_{\text{max}}=2$). Searching in the second time step will yield an expected reward of $\mu$. Settling for anything less than $\mu$ after the first time step would therefore result in a lesser expected reward in the second time step.
@@
We also prove that the optimal threshold ($T$) increases in the search time ($t_{\text{max}}$), and is finite as long as the search time is finite (Appendix A). Again, both results are intuitive: more search time increases the likelihood and cumulative payoff of achieving an ambitious target.
III.2 Overshooting the optimal threshold is costlier than undershooting
@@
An agent overshooting the optimal threshold is expected to receive lower rewards than an agent undershooting the optimal threshold by the same amount (Figs. 2a, 3, 4, and 5). This implies that uncertainty over the reward distribution should reduce the optimal satisfaction threshold (i.e., it motivates caution), analogous to the way that left-skewed uncertainty motivates risk aversion in other contexts [34]. However, in settings where agents learn about the reward distribution through searching and they can adjust their satisfaction thresholds accordingly, Kohn and Shavell [15] show that uncertainty can increase the optimal satisfaction threshold due to the information value of searching.
@@
Fig. 2b and c illustrate two real-world contexts in which people seem to behave as though they understand the search tradeoff that our theory captures. Fig. 2b compares the frequencies with which heterosexual men and women in four U.S. cities sent messages to prospective partners on a dating app, as a function of their prospective partners’ desirability compared to their own desirability (data from ref. [31]). Long-term partners tend to be similar in their desirability [6]. Thus, a zero desirability gap might be analogous to the mean reward ($\mu$) in our model. In the online dating market, both men and women expend the greatest (modal) search effort on prospective partners slightly more desirable than themselves, as our theory would predict (Fig. 2b).
@@
Fig. 2c shows the distribution of college applications from the 2008 U.S. high-school graduating class, comparing students’ own scores on the Scholastic Aptitude Test (SAT) to those of the median current student at the college they were applying to (data from ref. [32]). Attending a college where one has the median SAT score might be analogous to the mean reward ($\mu$) in our model. As in the dating example, the modal search (application) effort occurred near this mean college.
However, unlike the dating example, students sent more applications to weaker colleges than they sent to stronger ones. Hoxby [32] showed that most applications in this left tail came from low- and middle-income students, who faced greater non-academic constraints (e.g., geographic, economic) on which colleges they could attend. This low-ambition tail in college applications may also relate to the ‘safety school’ strategy, which prefers unambitious options in college to striking out and being left with no options. College applicants are able to pursue multiple strategies (applications) simultaneously (albeit with some costs from time and application fees), unlike the agents in our model.
@@
III.3 Reward landscape left skewness and ruggedness increase optimal satisfaction thresholds
For a landscape with smoothness $\varphi$, the variance of the associated AR(1) process is $\text{var}[X_{t}]=(1+\varphi)/(1-\varphi)$. The analytical expression for the expected cumulative reward on smooth landscapes can therefore be derived by simply scaling the threshold $T$ by a factor of $1/\sqrt{\text{var}[X_{t}]}$ in equation (2).
@@ III.3 Reward landscape left skewness and ruggedness increase optimal satisfaction thresholds
Rugged landscapes create higher-variance rewards, which increases achievable expected rewards and optimal satisfaction thresholds (Fig. 3a). Thus, agents should be more ambitious on rugged landscapes. The reward distribution for the first time period is the same for smooth and rugged landscapes, but autocorrelation reduces the subsequent variance on smooth landscapes. This means that agents can only make incremental changes: setting their sights too high will likely result in never encountering the desired reward, which would require a gradual climb due to the aurocorrelation.
@@
Figure 3: Effects of smoothness and skewness. Smoother reward landscapes yield lower cumulative rewards than the maximally rugged, non-skewed landscapes (a). Left-skewed reward landscapes yield higher cumulative rewards (b). Smooth landscapes are generated analytically. For the skewed landscape simulations, $\epsilon_{t}\sim\mathcal{SN}(0,1,\alpha)$, $\varphi=0$, $t_{\text{max}}=1000$, where the parameter $\alpha$ is varied to result in skew values from 0.8-0.8 to $0.8$, each averaged over $10^{4}$ simulations.
III.4 Social comparison hurts performance and penalizes ambition
We have previously assumed that agents know the reward distribution. What if they instead estimate the distribution, and evaluate their satisfaction, by looking at peers’ rewards [35]? This type of behavior is called social comparison [36, 37, 38, 39]. Indeed, in reality people often focus specifically on peers receiving higher rewards than they are—“upward social comparison” [29, 30].